OpenStudy (anonymous):

can someone show me the steps to this problem 17 + 2n = 21 - 2n

OpenStudy (anonymous):

first get the constants and the varibles on different sides by subtracting 17 from bothe sides and adding 2n to both sides...4n=4..then divide each side by 4...x=1

OpenStudy (saifoo.khan):

17 + 2n = 21 - 2n 2n + 2n = 21 - 17 4n = 4 n = 4/4 n = 1

OpenStudy (anonymous):

so is this a solution